### v1

@@ Abstract
This study provides a comparative analysis of state-of-the-art large language models (LLMs), analyzing how likely they generate vulnerabilities when writing simple C programs using a neutral zero-shot prompt. We address a significant gap in the literature concerning the security properties of code produced by these models without specific directives. N. Tihanyi et al. introduced the FormAI dataset at PROMISE ’23, containing 112,000 GPT-3.5-generated C programs, with over 51.24% identified as vulnerable. We expand that work by introducing the FormAI-v2 dataset comprising 265,000 compilable C programs generated using various LLMs, including robust models such as Google’s GEMINI-pro, OpenAI’s GPT-4, and TII’s 180 billion-parameter Falcon, to Meta’s specialized 13 billion-parameter CodeLLama2 and various other compact models. Each program in the dataset is labelled based on the vulnerabilities detected in its source code through formal verification using the Efficient SMT-based Context-Bounded Model Checker (ESBMC). This technique eliminates false positives by delivering a counterexample and ensures the exclusion of false negatives by completing the verification process. Our study reveals that at least 63.47% of the generated programs are vulnerable. The differences between the models are minor, as they all display similar coding errors with slight variations. Our research highlights that while LLMs offer promising capabilities for code generation, deploying their output in a production environment requires risk assessment and validation.
@@
Large Language Models, Vulnerability Classification, Formal Verification, Software Security, Artificial Intelligence, Dataset
@@ 1 Introduction
The introduction of Large Language Models (LLMs) is transforming software development and programming [1, 2, 3]. Every day, developers and computer scientists utilize various code creation and completion models to tackle different tasks [4, 5]. Research related to program synthesis using Generative Pre-trained Transformers (GPT) [6] is gaining significant traction, where initial studies indicate that the GPT models can generate syntactically correct yet vulnerable code [7]. A recent study conducted at Stanford University suggests that software engineers assisted by OpenAI’s codex-davinci-002 model were at a higher risk of introducing security flaws into their code [8]. As the usage of AI-based tools in coding continues to expand, understanding their potential to introduce software vulnerabilities becomes increasingly important. Given that LLMs are trained on data freely available on the internet, including potentially vulnerable code, there is a high risk that AI tools could replicate the same patterns. This leads to the question: Is it safe to employ these models in real software projects?
@@
As a first step towards answering this question, N. Tihanyi et al. published the FormAI dataset [9] at the 19th International Conference on Predictive Models and Data Analytics in Software Engineering (PROMISE’23). This dataset stands out as the first and largest collection of AI-generated compilable C programs with vulnerability classification, featuring $112,000$ samples. To guarantee the diversity of generated C codes, the authors developed a framework designed to produce a variety of programs that cover multiple coding scenarios, efficiently facilitating real-world bugs. The study employed Bounded Model Checking (BMC), a technique within Formal Verification (FV), to evaluate the security properties of the dataset. It revealed that $51.24$% of the C programs generated by GPT-3.5 were vulnerable. The original research discussed in [9] had three primary limitations that this comprehensive study aims to address:
  1. a.
    Firstly, it exclusively focused on OpenAI’s GPT-3.5 without evaluating other models. To bridge this gap, our study has been extended to encompass eight advanced LLMs, such as Google’s Gemini-pro [10], TII’s Falcon-180B [11], or Meta’s CodeLLama [12]. The models can be seen in Figure 1.
  2. b.
@@
    Secondly, the initial findings ($51.24$%) may have been under-reported due to the inherent limitations of BMC, indicating that the actual percentage of vulnerabilities could be higher. To address this issue, we transitioned our verification approach from bounded to unbounded verification, thereby enhancing the depth and accuracy of our security assessments [13, 14, 15, 16].
  3. c.
    Lastly, the complexity of the programs generated by GPT-3.5 was not thoroughly analyzed. To address these gaps, we have calculated the cyclomatic complexity CC [17] of each program generated by these LLMs to understand their intrinsic complexity better.
@@
Figure 1: The eight LLM models used in this study for code security comparison.
This is the first comprehensive study to compare vulnerabilities in code generated by various state-of-the-art LLMs. To summarize, this paper holds the following original contributions:
  • We present the FormAI-v2 dataset, consisting of $265,000$ compilable C program samples generated by eight different LLM models. Each C sample has been systematically labeled based on vulnerabilities identified through FV methods, particularly using the Efficient SMT-based Bounded Model Checker (ESBMC) [14, 15, 16] tool with an unbounded setting;
  • A detailed study to determine which LLMs produce code with the highest and the lowest number of vulnerabilities;
  • We offer a thorough analysis of the generated programs, including distributions of C keywords and the cyclomatic complexity. This analysis enables the comparison of LLMs in terms of coding complexity;
@@
The remaining sections of this paper are organized as follows: Section 2 provides an in-depth discussion on the motivation behind our research, outlining the primary questions that drove our study. Section 3 presents a comprehensive overview of the literature related, highlighting significant previous studies and their findings. Section 4 introduces the concepts of formal verification, focusing on the ESBMC module. Section 5 details the methodology we adopted to develop and label our dataset. Section 6 thoroughly evaluates our research findings and discusses their implications. Section 7 explores the limitations and threats to the validity of our study and proposes potential future research directions that could expand upon our results. Finally, Section 8 concludes the paper by summarising our contributions and addressing the research questions posed in this study.
@@ 2 Motivation
LLMs are typically employed for simpler tasks, such as crafting a prime number generation function or composing a basic program to sort an array, rather than tackling extensive projects involving thousands of lines of code [8]. The latest generation of LLMs can easily solve these simple tasks without facing any challenges. So far, the main area of interest in LLM-based code generation has been correctness. Datasets such as HumanEval [18] provide programming challenges to assess the performance of models in correctly solving various problems. For example, GPT-4 achieves a $67\%$ success rate in solving tasks compared to $48.1\%$ for GPT-3.5 [19]. On the contrary, even basic programs can pose challenges for the state-of-the-art LLM models regarding secure coding. To illustrate the issue, imagine a situation where a programmer asks GPT-4 the following: “Create a C program that prompts the user to input two numbers and then calculate their sum”. The code generated by GPT-4 is presented on the left in Figure 2, while the output from the formal verification tool ESBMC 7.5 is shown on the right.
@@
Figure 2: Motivation example: GPT-4 produced code with security vulnerabilities, demonstrated through formal verification results
@@
This simple code contains three potential security vulnerabilities. It exhibits an integer overflow during the addition of the variables number1 and number2, as well as two buffer overflows through the scanf() functions that retrieve input from the user. In $32$-bit computing architectures, integers are commonly stored as $4$ bytes ($32$ bits), which results in a maximum integer value of $2147483647$, equivalent to $12^{31}-$ 1. If one attempts to add $2147483647+1$ using this small program, the result will be incorrect due to integer overflow. The incorrect result will be -2147483648 instead of the expected $2147483648$. The addition exceeds the maximum representable value for a signed 32-bit integer $12^{31}-$ 1, causing the integer to wrap around and become negative due to the two’s complement representation.
@@
When GPT-4 is requested to write a secure version of this code using the following prompt: “Create a C program that prompts the user to input two numbers and then calculates their sum. Be careful and avoid security vulnerabilities.“, it only attempts to fix entering non-integer inputs by adding the following code snippet (Fig. 3):
@@
Figure 3: Response to a generic prompt requesting a secure version.
@@
Even after requesting a “secure” zero-shot prompt, all three original issues remain unresolved. Despite the significant advancements from GPT-3.5—which exhibited the same issue [9]—to GPT-4, our motivational example indicates that GPT-4 continues to produce code with vulnerabilities. Even if specifically requested in the prompt to avoid integer overflow in the program, the issue persists (Fig. 4):
@@
Figure 4: Zero-shot prompt requesting a fix for integer overflow
@@ 3.1 LLMs in Software Engineering
In software engineering (SE), it is essential to guarantee three main aspects of the code: correctness, safety, and security of the programs created. Functionally correct code should yield the expected outcomes for each input it processes. Safety in code means constructing fail-safe systems, protecting against accidental or unexpected inputs that might produce logically correct but undesirable results. Software security involves fortifying the software against external threats and deliberate attacks [21]. In a comprehensive study, Anwar et al. [22] highlight important safety issues related to LLMs beyond SE, from the disruptive socioeconomic impacts and cybersecurity risks to ethical issues. Vassilka et al. [23] discuss the need for SE education to adapt to AI advancements and prepare future software engineers to effectively and ethically utilize these technologies in their careers.
@@
Frameworks and techniques for turning prompts into executable code for SE are rapidly emerging, but the main focus is often functional correctness, omitting important security aspects [26, 27, 28]. There has been an arms race between researchers to excel in benchmarks using zero or few-shot frameworks [29, 30], multi-agent frameworks [31], fine-tuned models [32], and various other methods. As AI models evolve, their problem-solving capabilities improve significantly. However, whether these advancements also enhance the safety and security properties of the code they generate remains largely unclear and under-researched.
In [33], Lin et al. assessed different software process models to evaluate how these models affect code correctness (Pass@1333This metric highlights the model’s ability to produce correct and functional code on its first try without any revisions or corrections.). They also assessed the code quality of the AI-generated code by running static code checkers to uncover code smells [^4] Code smells are patterns in code that hint at potential problems, making maintenance harder but not necessarily causing immediate errors. They suggest areas where the code may need to be refactored for better quality and reliability.. This work had an interesting finding: the proposed software process models improved the quality of the generated code by significantly reducing code smells compared to what GPT-3.5 outputs by itself. Code smells or bad coding practices will not outright introduce vulnerabilities. However, several small-scale studies point to the fact that LLMs negatively impact software development from a security perspective. In [34], the authors generated $21$ small programs in five different languages: C, C++, Python, HTML, and Java. Combining manual verification with GPT-based vulnerability detection, the study found that only $5$ of the $21$ generated programs were initially secure.
@@
The programs are generated through a dynamic zero-shot prompting method, and the generated programs are not modified by any AI system afterward. While the primary goal of our paper is to investigate and compare the secure coding abilities of different LLM models, these conditions make the FormAI-v2 dataset suitable for ML purposes. On the other hand, AI models were trained on human-generated content; thus, the vulnerabilities produced have roots in incorrect code created by humans. Yet, as discussed in the next section, existing datasets notoriously include synthetic data (different from AI-generated), which can be useful for benchmarking vulnerability scanners but has questionable value for training purposes [54].
@@ 3.2 Existing Databases for Vulnerable C Code
Table 1: Comparison of various datasets based on their labeling classifications.
Specs Datasets Big-Vul Draper SARD Juliet Devign REVEAL Diverse Vul FormAI FormAI v2
@@
Dataset size 189k 1,274k 101k 106k 28k 23k 379k 112k 150k
Vul. Snippets 100% 5.62% 100% 100% 46.05% 9.85% 7.02% 51.24% 61%
Multi. Vulns.
Compilable
@@
B: By Design Vulnerable, F: Formal Verification with ESBMC, M: Manual Labeling
@@
For example, Draper contains C and C++ code from the SATE IV Juliet Test Suite, Debian Linux distribution, and public Git repositories. Since the open-source functions from Debian and GitHub were not labeled, the authors used a suite of static analysis tools: Cppcheck and Flawfinder [56]. However, the paper does not mention if vulnerabilities were manually verified or if any confirmation has been performed to root out false positives. In [54], on top of creating DiverseVul, Chen et al. merged all datasets that were based on GitHub commits and removed duplicates, thus making the most comprehensive collection of GitHub commits containing vulnerable C and C++ code.
@@ 3.3 Vulnerability Scanning and Repair
Manual verification involves human-driven processes such as code reviews and manual testing. While these methods effectively catch complex errors that automated tools might miss, they are labor-intensive and not scalable to large codebases or frequent updates. Static analysis evaluates source code without executing it, using static symbolic execution, data flow analysis, and control flow analysis. Style checking enforces coding standards for better readability and maintainability. These methods collectively enhance software integrity. The drawbacks are that this method can miss vulnerabilities that manifest only during runtime interactions and often introduce false positive results. Dynamic analysis tests the software’s behavior during execution [68]. It includes fuzzing, automated testing, run-time verification, and profiling. This technique requires executable code and often significant setup to simulate different environments and may not cover all execution paths.
@@
Formal Verification (FV) uses mathematical proofs to verify the correctness of algorithms against their specifications. It is the most rigorous form of software verification and is used in applications where reliability is critical, such as aerospace and medical devices. However, FV can be time-consuming and requires specialized knowledge, limiting its widespread adoption [21]. Recent advancements include machine learning techniques, particularly LLMs, in various aspects of software verification [69]. LLMs can assist in automated code review by suggesting improvements, detecting vulnerabilities, generating test cases, fixing bugs, and creating documentation. Despite their potential [70, 71, 72, 73, 74, 75, 76], LLMs, on their, own face limitations such as a lack of understanding of code semantics and difficulty in handling highly domain-specific knowledge [77], and they depend heavily on the quality and variety of the training data. Using LLMs as part of a framework to complement other techniques is, however, a promising area of research [78, 7, 9, 79].
@@
An earlier work from $2022$ examined the ability of various LLMs to fix vulnerabilities, where the models showed promising results, especially when combined. Still, the authors noted that such tools are not ready to be used in a program repair framework, where further research is necessary to incorporate bug localization. They further highlighted challenges in the tool’s ability to generate functionally correct code [80]. While LLMs struggle with detection by themselves, in [7], the authors demonstrated that GPT-3.5 could efficiently fix errors if the output of the ESBMC verifier is provided. Program repair is another emerging area where the application of LLMs is showing real promise, where in addition to fine-tuning strategies, the combination of LLMs with other tools appears to be an effective method [81, 82, 83, 41, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97]. In [98], the authors call for innovations to enhance automated vulnerability repair, particularly for developing more extensive training datasets to optimize LLMs.
3.4 Overview on Cyclomatic Complexity
Diversity is vital in our dataset to ensure a comprehensive representation of real-life vulnerabilities, facilitating a fair comparison across different LLMs. We add a new metric in FormAI-v2 to measure code complexity: cyclomatic complexity (CC). This metric, introduced by McCabe [17], quantifies the maintainability and complexity of programs by measuring the number of linearly independent paths within a program’s control-flow graph (CFG) [99]. It’s calculated as $V(G)$ for a graph $G$ with $n$ vertices, $e$ edges, and $p$ connected components.
$V(G)=e-n+2p.$
This metric, derived from the CFG – a directed graph of basic program blocks linked by control paths – is a numerical indicator of a program’s structural complexity. While complexity in functions increases due to multiple decision points, such as if-statements and loops imply higher testing and maintenance demands, critics note that CC might not fully capture complexities in modern software that extensively uses polymorphism and multi-threading. As will be discussed in Section 6, our findings suggest that cyclomatic complexity (CC) can indicate whether a model is capable of generating well-maintained, well-structured code or if it tends to produce poorly designed programs with high cyclomatic complexity, where higher numbers imply worse outcomes and lower numbers are preferable. Cyclomatic complexity can provide insights into the potential difficulty of testing, maintaining, or troubleshooting a piece of code and an indication of the likelihood of the code producing errors. Additionally, incorporating CC as a new feature in the dataset could enhance classification accuracy when used during the machine learning training phase.
The next section will offer necessary insights into formal verification to clarify the methodology employed in developing the dataset.

### v2

@@ Abstract
This study compares state-of-the-art Large Language Models (LLMs) on their tendency to generate vulnerabilities when writing C programs using a neutral zero-shot prompt. Tihanyi et al. introduced the FormAI dataset at PROMISE ’23, featuring 112,000 C programs generated by GPT-3.5-turbo, with over 51.24% identified as vulnerable. We extended that research with a large-scale study involving 9 state-of-the-art models such as OpenAI’s GPT-4o-mini, Google’s Gemini Pro 1.0, TII’s 180 billion-parameter Falcon, Meta’s 13 billion-parameter Code Llama, and several other compact models. Additionally, we introduce the FormAI-v2 dataset, which comprises 331 000 compilable C programs generated by these LLMs. Each program in the dataset is labeled based on the vulnerabilities detected in its source code through formal verification, using the Efficient SMT-based Context-Bounded Model Checker (ESBMC). This technique minimizes false positives by providing a counterexample for the specific vulnerability and reduces false negatives by thoroughly completing the verification process. Our study reveals that at least 62.07% of the generated programs are vulnerable. The differences between the models are minor, as they all show similar coding errors with slight variations. Our research highlights that while LLMs offer promising capabilities for code generation, deploying their output in a production environment requires proper risk assessment and validation. Please cite this once published: https://doi.org/10.1007/s10664-024-10590-1.
@@
Large Language Models, Vulnerability Classification, Formal Verification, Software Security, Artificial Intelligence, Dataset.
@@ 1 Introduction
Large Language Models (LLMs) are transforming software development and programming [1, 2, 3]. Every day, developers and computer scientists utilize various code creation and completion models to tackle different tasks [4, 5]. Research related to program synthesis using Generative Pre-trained Transformers (GPT) [6] is gaining significant traction, where initial studies indicate that GPT models can generate syntactically correct yet vulnerable code [7].
A study conducted at Stanford University suggests that software engineers assisted by OpenAI’s codex-davinci-002 model were at a higher risk of introducing security flaws into their code [8]. As the usage of AI-based tools in coding continues to expand, understanding their potential to introduce software vulnerabilities becomes increasingly important. Given that LLMs are trained on data freely available on the internet, including potentially vulnerable code, there is a high risk that AI tools could replicate the same patterns. This raises a critical question: Is it safe to employ these models in real projects?
@@
As a first step towards answering this question, Tihanyi et al. published the FormAI dataset [9] at the 19th International Conference on Predictive Models and Data Analytics in Software Engineering (PROMISE’23). This dataset is the first and largest collection of AI-generated compilable C programs with vulnerability classification, featuring 112 000112000112\,000112 000 samples. To guarantee the diversity of generated C codes, the authors developed a framework designed to produce a variety of programs that cover multiple coding scenarios, efficiently facilitating real-world bugs. The study employed Bounded Model Checking (BMC), a technique within Formal Verification (FV), to evaluate the security properties of the dataset. This initial study revealed that at least $51.24$% of the C programs generated by GPT-3.5-turbo were vulnerable.
Continuing the original research presented in [9], we aim to expand the scope of the study by addressing the key limitations highlighted by the research community. We identified four main limitations that we intend to address from the original paper:
  1. 1.
    The first paper exclusively focuses on OpenAI’s GPT-3.5-turbo, without evaluating other models. To bridge this gap, this paper compares nine state-of-the-art LLMs in secure coding, such as Google’s Gemini Pro 1.0 [10], OpenAI’s GPT-4o-mini, TII’s Falcon-180B [11], and Meta’s Code LLama 13B [12]. In addition, we have expanded the original dataset from 112 000112000112\,000112 000 to 331 000331000331\,000331 000 samples, where incorporating C code generated by different LLMs also enhances diversity.
  2. 2.
@@
    The initial findings on the percentage of vulnerable samples in the dataset ($51.24$%) may have been under-reported due to the limitations of bounded model checking, indicating that the actual percentage of vulnerabilities could be higher. To address this issue, we transitioned our verification approach from bounded to unbounded verification, thereby enhancing the depth and accuracy of our security evaluation [13, 14, 15, 16].
  3. 3.
    We have incorporated new labels into the dataset to enhance its usability for a broader research community. While all necessary features can be extracted and reproduced directly from the provided source codes, we have enhanced the dataset’s comprehensiveness by calculating the cyclomatic complexity (CC) [17] for each program, adding source lines of code (SLOC), including the exact stack trace for counterexamples, and providing a code snippet entry that captures only the 5 lines before and after the vulnerability. These additional features are valuable for machine learning tasks to help models generalize the problem and identify vulnerabilities more effectively and conduct more detailed comparisons in various research contexts.
  4. 4.
    To enhance the dataset, we have removed all Type 1, Type 2, Type 3-1, and Type 3-2 (with 10% deviation threshold) clones using the NiCad (Automated Detection of Near-Miss Intentional Clones) [18] tool. We note, that removing Type 3-2 clones with a larger threshold is not our goal. Even minor changes can be significant and determine whether a vulnerability is present or absent, potentially introducing different security risks. Moreover, different representations of a vulnerability can help models better generalize during machine learning training.
@@
RQ1: How does the security of LLM-generated code differ across various models? RQ2: What are the most typical vulnerabilities introduced during C code generation by different LLMs using neutral zero-shot prompts?
1.1 Main contribution
To summarize, this paper holds the following original contributions:
  • We present the FormAI-v2 dataset, consisting of 331 000331000331\,000331 000 compilable C programs (310 531310531310\,531310 531 with the exclusion of any Type 1, Type 2, Type 3-1 and Type 3-2 clones) generated by nine different LLMs. Each C sample has been systematically labeled based on vulnerabilities identified through formal verification methods, particularly using the Efficient SMT-based Bounded Model Checker (ESBMC) [14, 15, 16] tool with an unbounded setting;
  • A detailed study to determine which models produce code with the highest and the lowest number of vulnerabilities;
  • We provide a comprehensive analysis of the generated programs, detailing the distribution of vulnerabilities and highlighting the most frequently encountered types;
@@ 1.1 Main contribution
The remaining sections are organized as follows: Section 2 provides an in-depth discussion on the motivation. Section 3 presents a comprehensive overview of the literature related, highlighting significant previous studies and their findings. Section 4 introduces the concepts of formal verification, focusing on the ESBMC module. Section 5 details the methodology we adopted to develop and label our dataset. Section 6 presents our findings and discusses their implications. Section 7 explores the limitations and threats to the validity, and proposes potential future research directions. Finally, Section 8 concludes the paper by summarising our contributions and addressing the research questions posed in this study.
@@ 2 Motivation
In program synthesis, LLMs are generally used for simple tasks like writing a prime number generator or a basic program to sort an array, rather than handling large-scale projects involving thousands of lines of code [8]. The latest generation of LLMs can easily solve these simple tasks without facing any challenges. So far, the main area of interest in LLM-based code generation has been correctness. Datasets such as HumanEval [19] provide programming challenges to assess the performance of models in correctly solving various problems. For example, GPT-4 achieves a $67\%$ success rate in solving tasks compared to $48.1\%$ for GPT-3.5-turbo [20]. On the contrary, even for basic programming tasks, state-of-the-art LLMs may adopt insecure coding practices. To illustrate the issue, imagine a situation where a programmer asks GPT-4 the following: “Create a C program that prompts the user to input two numbers and then calculate their sum”. The code generated by GPT-4 is presented on the left in Figure 1, while the output from the formal verification tool ESBMC 7.6.1 is shown on the right.
@@
Figure 1: Motivation example: GPT-4 produced code with security vulnerabilities, demonstrated through formal verification results.
@@
This simple code contains three potential security vulnerabilities. It exhibits an integer overflow during the addition of the variables number1 and number2, as well as two buffer overflows through the scanf() functions that retrieve input from the user. In $32$-bit computing architectures, integers are commonly stored as $4$ bytes ($32$ bits), which results in a maximum integer value of 2 147 483 64721474836472\,147\,483\,6472 147 483 647, equivalent to $12^{31}-$ 1. If one attempts to add 2 147 483 647+1214748364712\,147\,483\,647+12 147 483 647 + 1 using this small program, the result will be incorrect due to integer overflow. The incorrect result will be 2 147 483 6482147483648-2\,147\,483\,648- 2 147 483 648 instead of the expected 2 147 483 64821474836482\,147\,483\,6482 147 483 648. The addition exceeds the maximum representable value for a signed 32-bit integer $12^{31}-$ 1, causing the integer to wrap around and become negative due to the two’s complement representation.
@@
When GPT-4 is requested to write a secure version of this code using the following prompt: “Create a C program that prompts the user to input two numbers and then calculates their sum. Be careful and avoid security vulnerabilities.”, it only attempts to fix entering non-integer inputs by adding the following code snippet (Figure 2):
@@
Figure 2: GPT-4 generated code snippet response after requesting a secure version of the code in Figure 1.
@@
Even after requesting a “secure” zero-shot prompt, all three original issues remain unresolved. Despite the significant advancements from GPT-3.5-turbo—which exhibited the same issue [9]—to GPT-4, our motivational example indicates that GPT-4 continues to produce code with vulnerabilities. Even if specifically requested in the prompt to avoid integer overflow in the program, the issue persists (Figure 3).
@@
Figure 3: Zero-shot prompt requesting a fix for integer overflow (failing to do so).
@@ 3.1 LLMs in Software Engineering
In software engineering (SE), it is essential to ensure three main aspects of the code: correctness, safety, and security of the programs created. Functionally correct code should yield the expected outcomes for each input it processes. Code safety means constructing fail-safe systems, protecting against accidental or unexpected inputs that might produce logically correct but undesirable results. Software security involves fortifying the software against external threats and deliberate attacks [22]. In a comprehensive study, Anwar et al. [23] highlight important safety issues related to LLMs beyond SE, from the disruptive socioeconomic impacts and cybersecurity risks to ethical issues. Vassilka et al. [24] discuss the need for SE education to adapt to AI advancements and prepare future software engineers to effectively and ethically utilize these technologies in their careers.
@@
Frameworks and techniques for turning prompts into executable code for SE are rapidly emerging, but the main focus mostly often functional correctness, omitting important security aspects [27, 28, 29], or reliability [30, 31, 32, 33, 34]. There has been an arms race between researchers to excel in correctness benchmarks using zero or few-shot frameworks [35, 36], multi-agent frameworks [37], fine-tuned models [38], and various other methods. As AI models evolve, their problem-solving capabilities improve significantly. However, whether these advancements also enhance the safety and security properties of the code they generate remains largely unclear and under-researched.
In [39], Lin et al. assessed different software process models to evaluate how these models affect code correctness (Pass@1333This metric highlights the model’s ability to produce correct and functional code on its first try without any revisions or corrections.). They also assessed the code quality of the AI-generated code by running static code checkers to uncover code smells [^4] Code smells are patterns in code that hint at potential problems, making maintenance harder but not necessarily causing immediate errors. They suggest areas where the code may need to be refactored for better quality and reliability.. This work had an interesting finding: the proposed software process models improved the quality of the generated code by significantly reducing code smells compared to what GPT-3.5-turbo outputs by itself. Code smells or bad coding practices will not outright introduce vulnerabilities. However, several small-scale studies point to the fact that LLMs negatively impact software development from a security perspective. In [40], the authors generated $21$ small programs in five different languages: C, C++, Python, HTML, and Java. Combining manual verification with GPT-based vulnerability detection, the study found that only $5$ of the $21$ generated programs were initially secure.
@@
The programs are generated through a dynamic zero-shot prompting method, and the generated programs are not modified by any AI system afterward. While the primary goal of our paper is to investigate and compare the secure coding abilities of different LLMs, these conditions make the FormAI-v2 dataset suitable for ML purposes. On the other hand, AI models were trained on human-generated content; thus, the vulnerabilities produced have roots in incorrect code created by humans. Yet, as discussed in the next section, existing datasets notoriously include synthetic data (different from AI-generated), which can be useful for benchmarking vulnerability scanners but has questionable value for training purposes [60].
@@ 3.2 Existing Databases for Vulnerable C Code
Table 1: Comparison of Various C Code Datasets
Specs Dataset Big-Vul Draper SARD Juliet Devign REVEAL Diverse Vul FormAI FormAI-v2
@@
Dataset size 189k 1274k 101k 106k 28k 23k 379k 112k 331k
Vul. Snippets 100% 5.62% 100% 100% 46.05% 9.85% 7.02% 51.24% 62.07%
Multi. Vulns. x x x x x x
Compilable x x x x x x
@@
B: By Design Vulnerable, FV: Formal Verification with ESBMC, M: Manual Labeling
@@
For example, Draper contains C and C++ code from the SATE IV Juliet Test Suite, Debian Linux distribution, and public Git repositories. Since the open-source functions from Debian and GitHub were not labeled, the authors used a suite of static analysis tools: CPPcheck [69] and Flawfinder [62]. However, the paper does not mention if vulnerabilities were manually verified or if any confirmation has been performed to root out false positives. In [60], on top of creating DiverseVul, Chen et al. merged all datasets that were based on GitHub commits and removed duplicates, thus making the most comprehensive collection of GitHub commits containing vulnerable C and C++ code.
@@ 3.3 Vulnerability Scanning and Repair
Manual verification involves human-driven processes such as code reviews and manual testing. While these methods effectively catch complex errors that automated tools might miss, they are labor-intensive and not scalable to large codebases or frequent updates. Static analysis evaluates source code without executing it, using static symbolic execution, data flow analysis, and control flow analysis. Style checking enforces coding standards for better readability and maintainability.
These methods collectively enhance software integrity. The drawbacks are that this method can miss vulnerabilities that manifest only during runtime interactions and often introduce false positive results. Dynamic analysis tests the software’s behavior during execution [74]. It includes fuzzing, automated testing, run-time verification, and profiling. This technique requires executable code and often significant setup to simulate different environments and may not cover all execution paths.
@@
Formal Verification (FV) uses mathematical proofs to verify the correctness of algorithms against their specifications. It is the most rigorous form of software verification and is used in applications where reliability is critical, such as aerospace and medical devices. However, FV can be time-consuming and requires specialized knowledge, limiting its widespread adoption [22]. Recent advancements include machine learning techniques, particularly LLMs, in various aspects of software verification [75]. LLMs can assist in automated code review by suggesting improvements, detecting vulnerabilities, generating test cases, fixing bugs, and creating documentation. Despite their potential [76, 77, 78, 79, 80, 81, 82], LLMs, on their own face limitations such as a lack of understanding of code semantics and difficulty in handling highly domain-specific knowledge [83], and they depend heavily on the quality and variety of the training data. Using LLMs as part of a framework to complement other techniques is, however, a promising area of research [84, 7, 9, 85]. An earlier work from $2022$ examined the ability of various LLMs to fix vulnerabilities, where the models showed promising results, especially when combined. Still, the authors noted that such tools are not ready to be used in a program repair framework, where further research is necessary to incorporate bug localization. They further highlighted challenges in the tool’s ability to generate functionally correct code [86].
@@
While LLMs struggle with detection by themselves, in [7], the authors demonstrated that GPT-3.5-turbo could efficiently fix errors if the output of the ESBMC verifier is provided. Program repair is another emerging area where the application of LLMs is showing real promise, where in addition to fine-tuning strategies, the combination of LLMs with other tools appears to be an effective method [87, 88, 89, 47, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103]. In [104], the authors call for innovations to enhance automated vulnerability repair, particularly for developing more extensive training datasets to optimize LLMs.